ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 240, 2020

Adiprasito, K. and E. Nevo Rigidity with few locations  711

Albiac, F., J. L. Ansorena, M. Cúth and M. Doucha Lipschitz free p-spaces for 0 < p < 1

65

Ansorena, J. L. See F. Albiac Borichev, A., A. Kononova and M. Sodin Notes on the Szegő

minimum problem. I. Measures with deep zeroes

725

Borichev, A., A. Kononova and M. Sodin Notes on the Szegő minimum problem. II. Singular measures

745

Corrêa. W. H. G. Complex interpolation of families of Orlicz sequence spaces

603

Costantini, M. Lyapunov exponents, holomorphic flat bundles and de Rham moduli space

345

Cúth, M. See F. Albiac Cyr, V. and B. Kra Realizing ergodic properties in zero entropy

subshifts

119

Doucha, M. See F. Albiac Duchesne, B., R. Tucker-Drob and P. Wesolek A new lattice

invariant for lattices in totally disconnected locally compact groups

539

Firsching, M. The complete enumeration of 4-polytopes and 3-spheres with nine vertices

417

Gallet, M. and J. Schicho Counting projections of rational curves  503

Glauberman, G., R. Guralnick, J. Lynd and G. Navarro Centers of Sylow subgroups and automorphisms

253

Guo, V. J. W. and M. J. Schlosser A family of q-hypergeometric congruences modulo the fourth power of a cyclotomic polynomial

821

Guralnick, R. See G. Glauberman Jiang, C. and Q. Wang Fusion rules for ℤ_{2}-orbifolds of affine and

parafermion vertex operator algebras

837

Jin, M., G.-B. Ren and I. Sabadini Slice Dirac operator over octonions

315

Kao, L. Pressure metrics and Manhattan curves for Teichmüller spaces of punctured surfaces

567

Katsoulis, E. Product systems of C^{∗}-correspondences and Takai duality 223 Keller, N. and O. Klein A structure theorem for almost low-degree

functions on the slice

179

Klein, O. See N. Keller Kononova, A. See A. Borichev Kra, B. See V. Cyr Li, M., Y. Shi, S. Wang and X. Wang Measures of intermediate

entropies for star vector fields

791

Lyall, N. and Á. Magyar Distances and trees in dense subsets of ℤ^{d}  769

Lynd, J. See G. Glauberman Magyar, Á. See N. Lyall Müller, P. F. X. and M. Passenbrunner Almost everywhere

convergence of spline sequences

149

Navarro, G. See G. Glauberman Nevo, E. See K. Adiprasito Passenbrunner, M. See P. F. X. Müller Pène, F. and B. Saussol Spatio-temporal Poisson processes for

visits to small sets

625

Pinamonti, A. and G. Speight Universal differentiability sets in Carnot groups of arbitrarily high step

445

Ramsey, N. Invariants related to the tree property  275

Ren, G.-B. See M. Jin Sabadini, I. See M. Jin Saussol, B. See F. Pène Schicho, J. See M. Gallet Schlosser, M. J. See V. J. W. Guo Shelah, S. Quite free complicated abelian groups, pcf and

Black Boxes

1

Shelah, S. Retraction of “Baire property and axiom of choice”  443

Shi, Y. See M. Li Smilansky, Y. Uniform distribution of Kakutani partitions generated

by substitution schemes

667

Sodin, M. See A. Borichev Speight, G. See A. Pinamonti Technau, N. and M. Widmer Counting lattice points and weak

admissibility of a lattice and its dual

99

Tucker-Drob, R. See B. Duchesne Wang, Q. See C. Jiang Wang, S. See M. Li Wang, X. See M. Li Wesolek, P. See B. Duchesne

Whang, J. P. Nonlinear descent on moduli of local systems  935

Widmer, M. See N. Technau

Williams, J. Normal subgroups of powerful p-groups  267

Ye, R. and E. Zelingher Exterior square gamma factors for cuspidal representations of GL_{n}: finite field analogs and level-zero representations

889

Zelingher, E. See R. Ye